OpenStudy (neverendingcycle):

Which shows the equation in slope-intercept form? 8x + 4y = 48 1. y = –2x + 12 2. y = 2x – 12 3. y = -1/2x + 12 4. y = 1/2x + 12

pooja195 (pooja195):

Slope intercept= y=mx+b m=slope b=y intercept

OpenStudy (shishi990):

Y=mx+b

pooja195 (pooja195):

Which one matches this form?

pooja195 (pooja195):

lol all of them we need to turn it into slope intercept form 8x + 4y = 48 What would the first step be?

OpenStudy (neverendingcycle):

all of them except # 2

pooja195 (pooja195):

Right but we need to find the slope intercept form for the given equation 8x + 4y = 48 We start by subtracting 8x from BOTH sides 4y=48-8x Re-arrange them into slope intercept form 4y=-8x+48 divide all of the numbers by 4

pooja195 (pooja195):

Dividing will isolate the y

OpenStudy (neverendingcycle):

y = -2x + 12?

pooja195 (pooja195):

Correct good job! :-)

OpenStudy (neverendingcycle):

Thanks for your help!

pooja195 (pooja195):

yw ^_^